# The exact travelling wave solutions of a KPP equation

### v6

@@ II The ODE which doesn’t contain explicitly the independent variable
Now let us be more specific and consider
$F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (12)
where $n>0$ and $a\geq 0$ are some constants, so (6) will take the form
$u_{\xi\xi}+cu_{\xi}=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (13)
$pp_{u}+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (14)
In search of elementary solutions of (14) let us return to Eq. (10), and in accordance with the well-known in mathematics principle, stating that the more general the problem is, the easier it is to solve it, let us even generalize it to
@@
$F(U)=F^{\prime}(0)\left[u-A(u)\right]\left[1+\chi A^{\prime}(u)\right],$ (19)
where $A(u)$ is some function, $\chi$ is some numerical parameter, and $\gamma(p)=c=$ const. In fact, taking
$\Phi(u)=F^{\prime}(0)\left[A(u)-u\right]$ (20)
@@
$P^{2}F^{\prime}(0)\left[A^{\prime}(u)-1\right]+Pc=1+\chi A^{\prime}(u).$ (21)
We see that choosing $P$ and $c$ we can turn (21) into the identity.
Now let us return to Eq. (14) and take
$\Phi(u)=u\left(u^{n}-1\right),\hskip 28.45274pt\Psi(u)=\left(u^{n}+a\right).$ (22)
Equation (18) takes the form
@@
In the vicinity of $u=1$, Eq. (10) becomes asimptotically
@@
In the vicinity of $u=0$, Eq. (10) becomes asimptotically
$pp_{v}+cp=F^{\prime}(0)u.$ (43)
Equation (43) has two solutions
@@
where $s_{3,4}$ are given by (36). Because both $s_{3}$ and $s_{4}$ have negative real part, both solutions will satisfy the boundary condition (7b). So we understood that the second boundary condition (11b) can be discarded – it is satisfied automatically.
@@
Our analysis of (29) allows us to understand (partially) the analytic properties of $p(u)$. If $c^{2}\geq 4F^{\prime}(0)$, from (37) and (38) we can expect that $p(u)$ would be either a single - or a double-valued function. (Further analysis is necessary to decide which option is realised.) Note that $c$ given by (26) obviously fulfills the inequality $c^{2}>4a$ for any $n>0$ and $a\geq 0$.)
@@ Appendix C Perturbation theory I
$\frac{du}{d\xi}=\frac{1}{c}F(u)\left(1-\frac{1}{c}\frac{dF(u)}{du}\right)$ (62)
@@
$\frac{d\xi}{c}=\frac{du}{F(u)}+\frac{1}{c}\frac{dF(u)}{F(u)}.$ (63)
@@
$\frac{\xi}{c}=\int\frac{du}{F(u)}+\frac{1}{c}\ln|F(u)|.$ (64)
@@ Appendix D Perturbation theory II
In Section II we have solved Eq. (13) for $c=c_{0}$. Motivated by the perturbation theory of the previous Section, let us study the solution of (13) for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series. To make life simpler $n=1$, we’ll consider only the case $n=1$. Presenting the solution of (14) (for $n=1$) as
$p(u)=p_{0}(u)+\frac{\delta c}{\sqrt{2}}\delta p(u),$ (70)
for $\delta p(u)$ in the lowest with respect to $\delta c$ order we obtain equation
$u(u-1)\delta p_{u}+\left(2u-1+c_{0}\right)\delta p+u(u-1)=0.$ (71)
@@
$\delta p(u)=\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\left(C-\int\frac{\left(1-u\right)^{c_{0}+1}du}{u^{c_{0}-1}}\right).$ (74)
It will be convenient to choose the lower limit in the in integral in (74) to be equal to zero. After that, for $\delta p$ to remain finite when $u\to 0,1$ we should choose $C=0$, thus obtaining
$\delta p(u)=-\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\int_{0}^{u}\frac{\left(1-u\right)^{c_{0}+1}du}{u^{c_{0}-1}}.$ (75)
A pleasant surprise is that $\delta p(u)$, as given by (75), goes to zero when $u\to 0,1$. (We couldn’t be sure in that just looking at Eq. (71), because the coefficient before $\delta p_{u}$ in the equation goes to zero in the above mentioned limits, hence one can doubt the validity of the linear approximation with respect to $\delta c$ in the vicinity of $u=0$ and $u=1$. )

### v7

@@ II The ODE which doesn’t contain explicitly the independent variable
In search of elementary solutions of (10), in accordance with the well-known in mathematics principle, stating that the more general the problem is, the easier it is to solve it, let us generalize it to
@@
$F(U)=\left[u-A(u)\right]\left[\chi A^{\prime}(u)+a\right],$ (16)
where $A(u)$ is some function, $\chi,a$ are some numerical parameters, and $\gamma(p)=c=$ const. In fact, taking
$\Phi(u)=\left[A(u)-u\right]$ (17)
@@
$P^{2}\left[A^{\prime}(u)-1\right]+Pc=\chi A^{\prime}(u)+a.$ (18)
Choosing $P^{2}=\chi$ and $c=P+a/P$ we turn (18) into identity.
Now let us be more specific and consider
$F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (19)
where $n>0$ and $a\geq 0$ are some constants, so that (6) will take the form
$u_{\xi\xi}+cu_{\xi}=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (20)
$pp_{u}+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (21)
We can take
$\Phi(u)=u\left(u^{n}-1\right),\hskip 28.45274pt\Psi(u)=\left(u^{n}+a\right),$ (22)
after which (15) takes the form
@@
In the vicinity of $u=1$, Eq. (10) becomes asymptotically
@@
In the vicinity of $u=0$, Eq. (10) becomes asymptotically
$pp_{u}+cp=-F^{\prime}(0)u.$ (43)
Equation (43) has two solutions (we’ll consider only the case $c^{2}>4F^{\prime}(0)$)
@@
where $s_{3,4}$ are given by (36). Because both $s_{3}$ and $s_{4}$ are negative, both solutions will satisfy the boundary condition (7b). So we understood that the second boundary condition (11b) can be discarded – it is satisfied automatically.
Note that for the case considered explicitly in the previous Section ($F(u)$ being given by (19) and $c$ being given by (26), obviously fulfilling the inequality $c^{2}>4a$ for any $n>0$ and $a\geq 0$) we get
$s_{3}$ $=-\frac{1}{\sqrt{n+1}},$ (45a)
$s_{4}$ $=-a\sqrt{n+1}.$ (45b)
Looking at Eq. (24) we realise that the solution (44a) is relevant in this case.
@@
Our analysis of (29) allows us to understand (partially) the analytic properties of $p(u)$. If $c^{2}\geq 4F^{\prime}(0)$, from (37) and (38) we can expect that $p(u)$ would be either a single - or a double-valued function. (Further analysis is necessary to decide which option is realised.)
@@ Appendix C Perturbation theory I
$\frac{du}{d\xi}=-\frac{1}{c}F(u)\left(1+\frac{1}{c}\frac{dF(u)}{du}\right)$ (63)
@@
$\frac{d\xi}{c}=-\frac{du}{F(u)}+\frac{1}{c}\frac{dF(u)}{F(u)}.$ (64)
@@
$\frac{\xi}{c}=-\int\frac{du}{F(u)}+\frac{1}{c}\ln|F(u)|.$ (65)
@@ Appendix D Perturbation theory II
In Section II we have solved Eq. (21) for $c=c_{0}$. Let us now study the solution of the equation for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the linear with respect to $\delta c$ approximation. To make life simpler, we’ll consider only the case $n=1$. Presenting the solution (for $n=1$) as
$p(u)=p_{0}(u)+\frac{\delta c}{\sqrt{2}}p_{1}(u)$ (71)
and substituting (71) into (21), for $p_{1}$ in linear approximation we obtain equation
$u(u-1){p_{1}}_{u}+\left(2u-1+c_{0}\right)p_{1}+u(u-1)=0.$ (72)
@@
$p_{1}(u)=\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\left(C-\int\frac{\left(1-u\right)^{c_{0}+1}du}{u^{c_{0}-1}}\right).$ (75)
Taking into account the boundary condition $p_{1}(1)=0$ we obtain
$p_{1}(u)=\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\int_{u}^{1}\frac{\left(1-v\right)^{c_{0}+1}dv}{v^{c_{0}-1}}.$ (76)
A question of the validity of the linear approximation presented above is quite delicate, because the coefficient before ${p_{1}}_{u}$ in Eq. (72) goes to zero when $u\to 0,1$. A good sign is that $p_{1}$ given by (76) has the asymptotic behavior $p_{1}\sim(u-1)$ in the vicinity of $u=1$, in accordance with (42). However, in the vicinity of $u=0$, $p_{1}$ has the asymptotic behavior $p_{1}\sim u$, as it should be in accordance with (44), only for $c_{0}>2$. (For $c_{0}<2$, $p_{1}\sim u^{c_{0}-1}$.) So we would like to believe that for $c_{0}>2$ the linear approximation is valid everywhere (for $0\leq u\leq 1$), and for $c_{0}<2$ the linear approximation is valid everywhere apart from the vicinity of $u=0$.